OpenStudy (anonymous):

?Help? Solve using the quadratic formula. 2/w^2 -11/w +9=0

OpenStudy (anonymous):

is it \[\frac{ 2 }{ w^2 } - \frac{ 11 }{ w } + 9 = 0 ??????\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

multiply whole equation by w^2, it will become quadratic equation. And you will be able to solve it.

OpenStudy (anonymous):

i did that and i got the wrong answer

OpenStudy (anonymous):

9w^2-11w+2=0 9w^2-9w-2w+2=0 (9w-2)(w-1)=0 w=2/9 and 1.

OpenStudy (anonymous):

ok thanks :) i see where i went wrong